# A Survey of 3D Reconstruction with Event Cameras

### v2

@@ VI Datasets
Outdoor navigation datasets: This category comprises ego-centric navigation sets, typified by DSEC [93][83], MVSEC [80], TUM-VIE [124], and ViViD++ [186]. These datasets align stereo events with LiDAR or tightly fused visual–inertial trajectories, yielding metre-scale ground truth that is indispensable for depth estimation and visual–inertial SLAM in outdoor environments. For urban driving scenarios, DSEC remains the reference standard, whereas TUM-VIE offers higher-resolution sensors and sub-centimetre inertial poses for precision studies.
@@
Indoor object and human reconstruction: This group targets object- and human-centric reconstruction in controlled indoor studios. EVIMO-2 [164] and MOEC-3D [106] provide millimetre-accurate meshes or per-pixel depth obtained from multi-camera motion capture or laser scanning, making them well suited to investigations of articulated or rigid-body shape recovery. DHP19 [185] is currently the sole public resource for full-body event capture, delivering four DAVIS346 views and millimetre-scale skeletons for non-rigid human modelling.
@@
Synthetic datasets: Where real scenes lack dense supervision, synthetic datasets provide idealised ground truth at scale. SynthEVox3D [95], SEVD [188], and EventScape [187] render voxel occupancies, semantic masks, and depth maps that enable large-scale pre-training and controlled ablation studies. While they offer reliable supervision for training, bridging the domain gap to real-world sparse inputs remains an open challenge, often addressed through augmentation strategies or adaptation techniques.
TABLE XI: Representative datasets used for event-based 3D Tasks. All the synthetic datasets are created by Blender [189]. Parts of the figures in the table are referenced from Ghosh et al. with permission [190].
@@
Dataset Sensors Frames Events Labels DSEC  [93][83] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] MVSEC  [80] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] TUM-VIE  [124] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] ViViD++  [186] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EV-IMO2  [164] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SynthEVox3D  [95] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EventScape  [187] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SEVD  [188] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
Photometric benchmarks: These datasets forego metric depth while delivering sharp RGB imagery with precise camera poses. EventNeRF [35] aligns asynchronous events with high-quality references for radiance-field research, while the classical DAVIS-240C sequences [101] continue to serve as a compact baseline for structure-from-motion and visual-odometry modules.
VII Research Gaps & Future Direction
@@
Standardised datasets and benchmarks: The lack of standardised datasets poses significant challenges. Existing datasets are either too limited, private, or inconsistent in terms of data formats, annotations, and scenarios. Consequently, researchers face difficulties in fair method comparison and reproducibility [26]. Future research should focus on developing comprehensive, publicly available datasets tailored explicitly for event-driven 3D reconstruction tasks. These datasets should cover diverse scenarios, including static and dynamic environments, challenging lighting conditions, and special object characteristics such as reflective or transparent surfaces. Additionally, establishing standardised evaluation protocols, including unified metrics for accuracy, robustness, and computational efficiency, will significantly aid in method validation and benchmarking.
@@
Non-frame-transferring synthetic event datasets: Currently, the main modelling platforms do not support event-based 3D modelling. While simulation tools such as ESIM [105] and Video-to-Event [191] exist, the event streams they generate rely on brightness changes between consecutive image frames, which introduces noticeable frame-based artefacts and fails to capture the sparse and asynchronous nature of real event data. The latest simulator for event cameras is showing the possibilities of creating such datasets [192, 193]. In the future, simulating event cameras within 3D modelling environments, which enables the collection of highly realistic event streams with perfect ground truth, holds great potential for deep learning in 3D reconstruction tasks.
Event representation: Event representation can be explored. Event representation significantly affects feature extraction, yet its impact on 3D reconstruction quality and efficiency remains underexplored. Although many novel representations have been proposed [194, 195, 27], no systematic comparison exists within a unified pipeline. Future research should rigorously evaluate these representations within unified frameworks to determine their strengths and weaknesses in different contexts. Further innovations could involve adaptive or hybrid event representations that dynamically adjust based on scene dynamics, lighting variations, and task-specific requirements, potentially leading to significant performance improvements. In addition, the development of event representation methods for feature enhancement for event-based 3D reconstruction tasks is also a possible direction.
Real-time reconstruction of dynamic scenes: Most existing event-driven 3D reconstruction methods primarily focus on static scenes. Dynamic environments introduce significant complexity, while dynamic rendering cannot achieve real-time. Future neural-rendering-based research should explore explicit modelling of temporal dynamics, incorporating deformable Neural Radiance Fields or dynamic 3D Gaussian Splatting with temporal encoding mechanisms. Moreover, real-time reconstruction of dynamic scenes holds great significance for applications in robotics, augmented reality, and autonomous driving, and is expected to become a key research direction. With real-time considerations, the method should focus more on the reconstruction of spatial occupation than on the reconstruction of features such as texture and colour.
@@
Experiments under extreme scenarios: It is well known that event cameras perform exceptionally under extreme conditions such as high-speed motion, low illumination, and high dynamic range. However, in the field of 3D reconstruction, comprehensive benchmarking against traditional cameras under such conditions remains limited [27]. Future work should design experiments to systematically evaluate and enhance the robustness of event cameras in these challenging scenarios. This may involve developing novel algorithms or sensor fusion strategies to achieve reliable reconstruction in environments that are difficult for conventional visual sensors.
@@
Reconstruction of object with challenging materials: Reconstructing 3D scenes with low-texture or non-Lambertian surfaces [196, 197] (e.g., glass or mirrors) remains highly challenging and is rarely studied. While the event camera’s sensitivity to photometric changes could potentially complement traditional cameras in such tasks, it may also be adversely affected. To date, only a few studies have explored using event cameras for photometric stereo of these types of objects [198]. In addition, dynamic and non-rigid objects introduce further complexity [199], as their continuously changing geometry violates the assumptions of many static reconstruction methods. Recent works in event-based dynamic NeRF [125, 153] show potential in this direction, but general solutions remain underexplored. Future research should investigate the effectiveness of event cameras in this context and, where applicable, design specialised algorithms tailored to reconstruct these challenging materials.
@@
Hardware and synchronisation constraints: The performance of event-driven pipelines is often bounded by hardware limitations, including the event camera’s resolution, timestamp precision, and bandwidth. In multimodal systems, accurate synchronisation between events, frames, IMU, and structured light sources is difficult to achieve and critical for fusion-based reconstruction [200]. Even small pose errors can propagate and degrade 3D structure estimation [71].
@@
Efficiency and scalability bottlenecks: NeRF and 3DGS methods offer high-fidelity reconstructions but are computationally expensive and memory-intensive. Real-time or large-scale scene reconstruction remains impractical, particularly in mobile or embedded scenarios. Efforts such as Event3DGS [167] and EVI-SAM [28] have started to address this, but further work is required on efficient architectures, pruning strategies, and event-specific network designs.
Underexplored modalities: While structured light and RGB-D sensors have been successfully combined with event cameras [7], modalities like LiDAR, polarisation, or event-based ToF remain underexplored. Their integration could provide additional priors to improve reconstruction under adverse conditions. Similarly, spatio-temporal fusion techniques offer promising opportunities but require careful synchronisation and encoding strategies [149].
@@
Exploring broader downstream applications: Due to the limitations of existing datasets, current research on event-based 3D reconstruction remains constrained to scenarios supported by available benchmarks. However, there are many foreseeable downstream applications yet to be fully explored:
@@ VIII Conclusion
Chuanzhi Xu proposed the review topic, collected most of the referenced papers, and conducted and structured the majority of the work. He also created the majority of the tables and figures. Haoxian Zhou contributed significantly to the writing of several sections and produced multiple figures. Langyi Chen wrote Section 6 on datasets. Haodong Chen and Qiang Qu carefully reviewed the manuscript, and Qiang Qu additionally contributed by revising the paper, coordinating the project, and managing subsequent communications. Ying Zhou, Vera Chung, and Weidong Cai supervised the project, provided necessary materials, and offered general guidance and feedback on the manuscript.

### v3

@@ 6.1 Outdoor navigation datasets
This category comprises ego-centric navigation sets, typified by DSEC [[S7-13]][DSEC2], MVSEC [MVSEC], TUM-VIE [TUM-VIE], and ViViD++ [ViViD++]. These datasets align stereo events with LiDAR or tightly fused visual–inertial trajectories, yielding metre-scale ground truth that is indispensable for depth estimation and visual–inertial SLAM in outdoor environments. For urban driving scenarios, DSEC remains the reference standard, whereas TUM-VIE offers higher-resolution sensors and sub-centimetre inertial poses for precision studies.
6.2 Indoor object and human reconstruction
@@
This group targets object- and human-centric reconstruction in controlled indoor studios. EVIMO-2 [EVIMO2] and MOEC-3D [[MD5]] provide millimetre-accurate meshes or per-pixel depth obtained from multi-camera motion capture or laser scanning, making them well suited to investigations of articulated or rigid-body shape recovery. DHP19 [DHP19] is currently the sole public resource for full-body event capture, delivering four DAVIS346 views and millimetre-scale skeletons for non-rigid human modelling.
6.3 Synthetic datasets
@@
Where real scenes lack dense supervision, synthetic datasets provide idealised ground truth at scale. SynthEVox3D [[Dense]], SEVD [SEVD], and EventScape [EventScape] render voxel occupancies, semantic masks, and depth maps that enable large-scale pre-training and controlled ablation studies. While they offer reliable supervision for training, bridging the domain gap to real-world sparse inputs remains an open challenge, often addressed through augmentation strategies or adaptation techniques.
Table 11: Representative datasets used for event-based 3D Tasks.
@@
Dataset Sensors Frames Events Labels DSEC  [[S7-13]][DSEC2] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] MVSEC  [MVSEC] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] TUM-VIE  [TUM-VIE] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] ViViD++  [ViViD++] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EV-IMO2  [EVIMO2] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SynthEVox3D  [[Dense]] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] EventScape  [EventScape] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] SEVD  [SEVD] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
6.4 Photometric benchmarks
@@
These datasets forego metric depth while delivering sharp RGB imagery with precise camera poses. EventNeRF [[N3]] aligns asynchronous events with high-quality references for radiance-field research, while the classical DAVIS-240C sequences [DAVIS240C] continue to serve as a compact baseline for structure-from-motion and visual-odometry modules.
7 Metrics
Appropriate evaluation metrics are essential for fair comparison and objective analysis of event-based 3D reconstruction methods. Since different approaches vary in reconstruction targets and output representations, no single metric can fully characterize reconstruction quality. As a result, existing works adopt evaluation criteria that are tailored to specific method categories and output forms. In this section, we summarise the commonly used evaluation metrics and organize them according to different classes of event-based 3D reconstruction methods.
For traditional stereo and monocular geometry-based methods, multimodal methods with traditional outputs, the primary focus lies on the geometric accuracy of the reconstructed point clouds or meshes. Metrics such as Mean Error, Median Error, Relative Error, Root Mean Square Error (RMSE), and Mean Absolute Error (MAE) are employed to quantify the deviation in spatial occupancy and geometric consistency between the reconstruction results and the ground truth [[S0], [S2], [S4], [XS2], [S6], [S7], [S8], [XS3], [XS4], [EMVSconf], [M3], [M4], [XDMT3], [XDMT1], [DMT2], [XDMT2], [DMT4], [DMT5], [DMT6]]. For monocular learning-based methods, Intersection over Union (IoU) and F-Score are adopted as the principal metrics to measure the consistency of object shape recovery [[MD1], [Dense], [Towards]]. IoU serves as the most intuitive indicator for measuring the degree of overlap between two volumes (the reconstructed volume and the ground truth volume), assessing the consistency of spatial occupancy between the object shape recovered by the algorithm and the real object. The F-score is the harmonic mean of precision and recall; it requires not only accuracy in the reconstructed points but also completeness in the reconstruction, thereby providing a better comprehensive evaluation of the recovery effects regarding object boundaries and details.
Regarding neural rendering and novel view synthesis methods such as NeRF and 3DGS [[N3], [N2], [XN2], [G10], [G5], [G8], [N3], [N14], [XG4], [XG6]], the emphasis is placed on assessing the photometric fidelity and perceptual quality of the rendered images. Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) are utilized as the core evaluation standards. PSNR reflects the purity of the image at the signal level, where a higher PSNR value indicates smaller pixel value errors and less distortion. SSIM attempts to simulate the human eye’s perception capability regarding image structure. Unlike PSNR, which focuses solely on pixel value differences, SSIM compares two images across three dimensions: luminance, contrast, and structure. It posits that the human eye is more adept at capturing structural information within an image, such as object contours and texture orientation. LPIPS leverages pre-trained deep neural networks (simulating the human visual cortex) to extract image features and calculates the distance between these features; a lower LPIPS value signifies that the rendered image appears more natural and realistic to the human eye.
8 Research Gaps & Future Direction
@@
8.1 Standardised datasets and benchmarks
Despite the rapid methodological progress in event-based 3D reconstruction, the field continues to suffer from a scarcity of large-scale, real-world, and openly accessible datasets explicitly designed for reconstruction tasks [[R2], [R3]].
Many existing works rely on private datasets or synthetic event generation pipelines [[XS2], [M2], [XDMT3], [DMT4], [N12], [G1]], which significantly hinders fair comparison, reproducibility, and long-term benchmarking across methods. Publicly available datasets are often limited in scene diversity, confined to controlled laboratory environments, or released without dense geometric supervision, referring to Section 6.
To address this bottleneck, several community-level initiatives and collaborative efforts are essential. First, standardised data acquisition pipelines that integrate event cameras with mature sensing modalities—such as LiDAR, RGB-D sensors, and motion-capture systems (e.g., Vicon) - should be promoted to reduce the technical and logistical barriers to dataset construction while ensuring high-quality geometric ground truth. Second, multi-institution collaborations, similar to those established in autonomous driving and robotics research, would enable the collection of geographically and scenically diverse datasets under unified calibration, annotation, and synchronization protocols. Third, the establishment of shared benchmarks, challenges, and public leaderboards, analogous to KITTI [Geiger2012KITTI] or TUM [Sturm2012TUM] in conventional vision, could provide strong incentives for open data release, sustained dataset maintenance, and transparent method comparison. Finally, physically grounded simulators and differentiable event-generation models offer a complementary pathway to augment real-world data [Hu2021V2E, [esim]], provided that domain gaps between synthetic and real events are systematically quantified and mitigated through hybrid real–synthetic evaluation protocols.
These initiatives are critical for transforming event-based 3D reconstruction from a collection of isolated, dataset-specific studies into a unified, reproducible, and scalable research ecosystem.
8.2 Synthetic event datasets not reliant on frame interpolation
@@
Currently, the main modelling platforms do not support event-based 3D modelling. While simulation tools such as ESIM [[esim]] and Video-to-Event [[v2e]] exist, the event streams they generate rely on brightness changes between consecutive image frames, which introduces noticeable frame-based artefacts and fails to capture the sparse and asynchronous nature of real event data.
The latest simulator for event cameras is showing the possibilities of creating such datasets [eventSimulator, eventSimulator2]. In the future, simulating event cameras within 3D modelling environments, which enables the collection of highly realistic event streams with perfect ground truth, holds great potential for deep learning in 3D reconstruction tasks.
8.3 Event representation
Event representation remains one of the most critical and promising research directions in event-based 3D reconstruction. Existing studies have shown that the choice of event representation has a substantial impact on feature extraction and learning behavior. However, its influence on reconstruction accuracy, robustness, and computational efficiency has not yet been systematically explored. Although a variety of event representations have been proposed in recent years [[Eventrep1], [Eventrep2], [Towards]], there is still a lack of comprehensive comparison under a unified reconstruction framework and evaluation protocol, which limits a clear understanding of their relative strengths and weaknesses.
Future research should first focus on conducting rigorous evaluations of representative event encodings within a unified 3D reconstruction pipeline. Such studies should analyze their performance in terms of geometric accuracy, temporal consistency, robustness, and computational cost, particularly across different scenarios such as static and dynamic scenes, high-speed motion, low-texture environments, and challenging illumination conditions. This would provide task-driven guidance for selecting appropriate event representations in practical 3D reconstruction systems. Beyond fixed representations, a particularly promising direction lies in adaptive and hybrid event representations. Instead of relying on predefined temporal windows or spatial aggregation strategies, these approaches can dynamically adjust their spatiotemporal encoding according to scene dynamics, event density, motion intensity, or task requirements. For example, high temporal resolution representations may be favored in rapidly moving regions, while stronger spatial aggregation can be applied to more stable structures, potentially achieving a better balance between reconstruction accuracy and computational efficiency.
In addition, developing geometry-aware event representations constitutes an important research avenue. Compared to generic event encodings, such representations can explicitly emphasize geometrically informative cues such as edges, surface discontinuities, and structural consistency, thereby providing more discriminative inputs for downstream tasks including depth estimation, volumetric reconstruction, and neural rendering.
8.4 Real-time reconstruction of dynamic scenes
Most existing event-driven 3D reconstruction methods primarily focus on static scenes, while extending them to dynamic environments introduces substantial challenges in both modelling and computation [[R1], [EMVS]]. In particular, achieving real-time performance alongside high-fidelity reconstruction remains challenging for neural-rendering-based approaches, especially in the presence of non-rigid motion and temporal variation [Pumarola2021DNeRF]. Rather than treating dynamic reconstruction as a direct extension of static pipelines, recent studies suggest that explicitly modelling temporal dynamics constitutes a more promising research direction.
Future research should therefore focus on representations and models that explicitly encode motion and temporal evolution, such as deformable Neural Radiance Fields or dynamic 3D Gaussian Splatting augmented with temporal embeddings or motion-aware primitives [Pumarola2021DNeRF].
From a real-time perspective, another promising avenue lies in prioritizing spatial occupancy and coarse geometry over photorealistic appearance. For many downstream applications such as robotics, augmented reality, and autonomous driving, accurate reconstruction of scene structure and free space is often more critical than detailed texture or colour reproduction [[EMVS], [M1]]. This observation motivates lightweight representations that focus on dynamic occupancy, depth, or surface geometry, while simplifying or deferring appearance modelling.
The most promising directions for real-time dynamic scene reconstruction are likely to emerge from the combination of motion-aware representations, event-driven temporal modelling, and task-oriented simplifications that balance reconstruction fidelity with computational efficiency.
8.5 Experiments under extreme scenarios
@@
It is well known that event cameras perform exceptionally under extreme conditions such as high-speed motion, low illumination, and high dynamic range. However, in the field of 3D reconstruction, comprehensive benchmarking against traditional cameras under such conditions remains limited [[Towards]]. Future work should design experiments to systematically evaluate and enhance the robustness of event cameras in these challenging scenarios. This may involve developing novel algorithms or sensor fusion strategies to achieve reliable reconstruction in environments that are difficult for conventional visual sensors.
8.6 Reconstruction of object with challenging materials
@@
Reconstructing 3D scenes with low-texture or non-Lambertian surfaces [zhang2004shape, li2020inverse] (e.g., glass or mirrors) remains highly challenging and is rarely studied. While the event camera’s sensitivity to photometric changes could potentially complement traditional cameras in such tasks, it may also be adversely affected. To date, only a few studies have explored using event cameras for photometric stereo of these types of objects [[EventPS]]. In addition, dynamic and non-rigid objects introduce further complexity [Non-Rigid], as their continuously changing geometry violates the assumptions of many static reconstruction methods. Recent works in event-based dynamic NeRF [[N4], [N14]] show potential in this direction, but general solutions remain underexplored. Future research should investigate the effectiveness of event cameras in this context and, where applicable, design specialised algorithms tailored to reconstruct these challenging materials.
8.7 Hardware and synchronisation constraints
@@
The performance of event-driven pipelines is often bounded by hardware limitations, including the event camera’s resolution, timestamp precision, and bandwidth. In multimodal systems, accurate synchronisation between events, frames, IMU, and structured light sources is difficult to achieve and critical for fusion-based reconstruction [10901942]. Even small pose errors can propagate and degrade 3D structure estimation [[S8]].
8.8 Efficiency and scalability bottlenecks
@@
NeRF and 3DGS methods offer high-fidelity reconstructions but are computationally expensive and memory-intensive. Real-time or large-scale scene reconstruction remains impractical, particularly in mobile or embedded scenarios. Efforts such as Event3DGS [[G4]] and EVI-SAM [[M3]] have started to address this, but further work is required on efficient architectures, pruning strategies, and event-specific network designs.
8.9 Underexplored modalities
While structured light and RGB-D sensors have been successfully combined with event cameras [li2022high], other modalities such as LiDAR, polarisation cameras, and event-based time-of-flight sensors remain comparatively underexplored. Their integration could introduce complementary geometric and physical priors, particularly under fast motion or challenging illumination.
In multimodal 3D reconstruction, effective fusion strategies focus on complementary sensing rather than simple data aggregation. Event cameras provide high-temporal-resolution motion cues but lack spatial density and absolute scale, whereas LiDAR and RGB-D sensors offer metric geometric constraints but are sensitive to motion artifacts, and IMU measurements ensure short-term motion consistency at the cost of long-term drift.
Promising approaches therefore perform fusion at both the representation and optimization levels. At the front end, accurate temporal synchronization enables events to capture rapid motion while depth or LiDAR measurements act as stable geometric anchors and IMU signals constrain pose continuity. At the back end, joint optimization or neural representation frameworks, such as multimodal NeRF or 3D Gaussian Splatting, map heterogeneous observations into a unified 3D representation with modality-specific losses and adaptive weighting to suppress sensor-dependent noise.
Importantly, selective modality activation based on scene dynamics represents a particularly effective strategy. For example, event data can be emphasized in high-speed or high-dynamic-range scenarios, while depth or LiDAR cues dominate in structurally stable regions, enabling improved reconstruction quality and real-time performance [[DMT5]].
8.10 Exploring broader downstream applications
@@
Due to the limitations of existing datasets, current research on event-based 3D reconstruction remains constrained to scenarios supported by available benchmarks. However, there are many foreseeable downstream applications yet to be fully explored:
@@ 9 Conclusion
Chuanzhi Xu proposed the review topic, collected most of the referenced papers, and conducted and structured the majority of the work. He also created the majority of the tables and figures. Haoxian Zhou contributed significantly to the writing of several sections and produced multiple figures. Langyi Chen wrote Section 6 on datasets. Haodong Chen, Zeke Zexi Hu, and Zhicheng Lu carefully reviewed and revised the paper. Qiang Qu coordinated the project and managed subsequent communications. Ying Zhou, Vera Chung, and Weidong Cai supervised the project, provided necessary materials, and offered general guidance and feedback on the manuscript.